OpenStudy (anonymous):

Exponential Functions: f(x)=3^x I need to find a way to graph this, but I am kind of lost. I made a "T Table" and inserted -1, -2, and 2 for x. for f(x) I got, respectively, -3, 9, and 9. Am I doing this right? And will this be a parabola?

OpenStudy (anonymous):

Hey L2theM...unfortunately, you're not doing it right... You have\[f(x)=3^x\]which is an exponential function. It isn't a parabola. Exponentials grow in the positive x-direction (i.e. go to infinity as x goes to infinity) and shrink to 0 as x goes to minus infinity. The function is NEVER 0 or negative.

OpenStudy (anonymous):

\[f(-1)=3^{-1}=\frac{1}{3^1}=\frac{1}{3}\]\[f(-2)=3^{-2}=\frac{1}{3^2}=\frac{1}{9}\]\[f(2)=3^2=9\]Do you know your power laws? Follow the pattern and you should be right.